OpenStudy (anonymous):

Find cos 2A, Sin A = -2/9

OpenStudy (anonymous):

You will need to use some identities here:\[\bf \cos(2a)=1-2\sin^2(a)\]We know what sin(a) is, so we just plug that in to the identity and evaluate.\[\bf \cos(2a)=1-2(-2/9)^2=..?\]Can you evaluate that? @walter9906

OpenStudy (anonymous):

Thank you. I have one more. find Cos 2A \[\cos A=\sqrt{3}/4\]